OpenStudy (horsegirl27):

Help! I'm so confused what I'm doing!! I will medal and fan. I will post the question below.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i hope it is not geometry !

OpenStudy (horsegirl27):

don't worry it's not!

OpenStudy (horsegirl27):

That's the first half. Explain what I need to do then please help me do it?

OpenStudy (horsegirl27):

I understood the lesson pretty well, so I don't think the actually doing it will be hard, I just am not sure what I'm doing and would like some help working on it

OpenStudy (horsegirl27):

@thomaster @TheEdwardsFamily @shrutipande9 @Somy @.Sam. @goatgal102

OpenStudy (theedwardsfamily):

@iambatman @phi @directrix @pooja195

OpenStudy (theedwardsfamily):

@Kainui @thomaster @.Sam.

OpenStudy (theedwardsfamily):

@undeadknight26 @k_lynn

OpenStudy (theedwardsfamily):

im trying to tag as many people that can help you. @horsegirl27

OpenStudy (theedwardsfamily):

:)

OpenStudy (horsegirl27):

Thank you. It feels like no one is on

OpenStudy (theedwardsfamily):

ikr

OpenStudy (anonymous):

@133 @HaileyD @MinightTheWolfie will somone just help this girl it's confusing for me 2

OpenStudy (anonymous):

@SamsungFanBoy

OpenStudy (horsegirl27):

@amorfide @shrutipande9 @Somy @sreekar369 @sjman1 @HaileyD @Catch.me

OpenStudy (anonymous):

Did you look for any company??

OpenStudy (anonymous):

@horsegirl27

OpenStudy (amorfide):

I can not do it, please don't tag me again

OpenStudy (horsegirl27):

@Catch.me I'm not really sure what I'm doing in the first step, so no and @amorfide sorry for tagging you again and thanks for telling me you can't help.

OpenStudy (tkhunny):

Why don't you stop tagging people and start showing your work. You CAN'T have no idea. What are your initial thoughts?

OpenStudy (tkhunny):

It says "research". Why don't you do that? http://www.interest.com/cd-rates/news/2-year-cd-rates/ Think you can find information about 5-year rates?

OpenStudy (horsegirl27):

Yeah let me look

OpenStudy (horsegirl27):

@tkhunny https://www.depositaccounts.com/cd/5-year-cd-rates.html?gclid=Cj0KEQjwuI-oBRCEi87g0K3O8OoBEiQAb25WAaTmEpbezPC8OHzd1frOp2fUR9Ys2Mtbwic5Os48oOYaAm4y8P8HAQ is this what I need? Now I have a lot of information that I don't know what to do with and which parts to use

OpenStudy (horsegirl27):

Then, do I find a company? See, idk what to do

OpenStudy (horsegirl27):

omg why is no one helping!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (horsegirl27):

@phi can you please help me?

OpenStudy (tkhunny):

Calm down and think. What is needed? Stop the panic. Stop the tagging. You are right that no one is helping. You should count yourself in that list of non-helpers. Seriously, panic down and think. Now you have a list of rates, 2-Year and 5-Year. What other requirements are stated? Find one with a $5,000 minimum. There are no other restrictions. Just pick one. Read the problem statement a couple more times. Calmly...

OpenStudy (horsegirl27):

ok

OpenStudy (horsegirl27):

So for 2 year and 5 year I found one each that have a minimum of 5000. What do I do next? Do I choose one? which one? Both because one is 2 year one is 5 year? See that's where I'm confused:choosing. Should I say here which ones I choose? ik I'm trying to be calm but just idk

OpenStudy (tkhunny):

Re-reading the problem statement, you can find any with a minimum as much as $5,000. So, you could pick one with no minimum or a $1,000 minimum, just not one with a $25,000 minimum. Follow the instructions. Fill in the chart. Write down the rate, both 2-, and 5-year Write down the company names. Make a few calculations, I guess assuming the 2-year rolls over twice. Just add $50 each year for Mr. "Investor".

OpenStudy (tkhunny):

Investor 5000 5050 5100 5150 5200 5250 Done with that one.

OpenStudy (horsegirl27):

Ok, thanks. I will do that and then could I post what I get here? Or should I continue if I understand?

OpenStudy (tkhunny):

Just show your work. If I selected a 5-year rate of 2.17%, I would do this: 5000 5000*1.0217 = $5,108.50 5000*1.0217^2 = $5,219.35 5000*1.0217^3 = $5,332.61 5000*1.0217^4 = $5,448.33 5000*1.0217^5 = $5,566.56 That's all. Re-read the problem statement. Make sure you know what it is asking. Take a couple deep breaths.

OpenStudy (horsegirl27):

Thank you!

OpenStudy (horsegirl27):

Oh, I actually understand now what I'm supposed to be doing. Thanks so much @tkhunny!!

OpenStudy (horsegirl27):

@tkhunny here's an update on what I have: 2 year company name: synchrony bank, minimum is $2000, the rate is 1.35% so my work is this: 2000 2000*1.0135=2027 2000*1.0135^2=2054.40 I will do it the same way for the 5-year company. Am I doing it the way I need to be doing it, or have I made a mistake?

OpenStudy (horsegirl27):

I'm not sure if I did the next part right. the minimum of the bank I chose was 500, and the rate was 2.27% and the final answer I got was 500*1.0227^5=572.10 Am I supposed to add that to 5000? Or am I right with the answer I have?

OpenStudy (horsegirl27):

@sammixboo Do you know?

Mehek (mehek14):

I did this for FLVS

OpenStudy (horsegirl27):

Ok

OpenStudy (horsegirl27):

Could you help me out with it?

OpenStudy (horsegirl27):

So far, am I doing it right?

Mehek (mehek14):

let me check

OpenStudy (phi):

they want you to put in $5000 for each of the CD's

OpenStudy (horsegirl27):

thanks. I'm just not sure about my work for the 5-year plan and questions 5-8

Mehek (mehek14):

for the first one you have to find a company

OpenStudy (horsegirl27):

I found companies and wrote down the information about them that I needed

OpenStudy (horsegirl27):

You can see what tk helped me with and what I did

OpenStudy (phi):

2 year company name: synchrony bank, minimum is $2000, the rate is 1.35% so my work is this: 2000 2000*1.0135=2027 2000*1.0135^2=2054.40 **** you have to put in 5000 (not 2000) the reason they asked for a minimum of at least 5000 is so you can put in 5000

Mehek (mehek14):

for the second one you have to make functions based on the companies you found

OpenStudy (horsegirl27):

oh ok

OpenStudy (horsegirl27):

so just put in 5000

OpenStudy (phi):

yes. put in 5000 into each CD

OpenStudy (horsegirl27):

ok

Mehek (mehek14):

the third one is a linear function

Mehek (mehek14):

that should be easy to make they give you the y-intercept and the rate of change

OpenStudy (horsegirl27):

so then it will be 5000*1.0135^2 and 5000*1.0227^5?

Mehek (mehek14):

and then for 4, you just use the functions to find how much money in the years

OpenStudy (horsegirl27):

And for the third it's what tk showed up there right^^^ 5000 5050 5100....

OpenStudy (horsegirl27):

Slow down @Mehek14 you're going a little too fast

OpenStudy (horsegirl27):

Wait till I understand before explaining the next one please

Mehek (mehek14):

\(\Large{f(x) = 5000(1.0135)^2}\) \(\Large{f(x) = 5000(1.0227)^2}\)

OpenStudy (horsegirl27):

Wait shouldn't the second be 5 not 2?

Mehek (mehek14):

my bad let me fix it

OpenStudy (horsegirl27):

Thanks

Mehek (mehek14):

\(\Large{f(x) = 5000(1.0227)^5}\)

OpenStudy (horsegirl27):

Ok. So those are for number 2

Mehek (mehek14):

now make the function for the investor's plan

OpenStudy (horsegirl27):

f(x)=50x+5000?

Mehek (mehek14):

actually, the functions are \(\Large{f(x) = 5000(1.0135)^x}\) \(\Large{f(x) = 5000(1.0227)^x}\) x = years

Mehek (mehek14):

so here are your three functions: \(\Large{f(x) = 5000(1.0135)^x}\) \(\Large{f(x) = 5000(1.0227)^x}\) \(\Large{f(x) = 50x + 5000}\)

OpenStudy (horsegirl27):

ok

Mehek (mehek14):

plug in the number of years for the table

OpenStudy (horsegirl27):

that answers 1-3

Mehek (mehek14):

yes for 4 you just substitute x with the number of years

OpenStudy (horsegirl27):

ok so for the table... let me do that

Mehek (mehek14):

you can do that on your own

OpenStudy (horsegirl27):

Wait yeah so for 4 I'm just finding the amount for each year. I will do that later

Mehek (mehek14):

so I'm done helping?

OpenStudy (horsegirl27):

Can I post the next questions and have you explain those? I found those a little harder

Mehek (mehek14):

If I know them

Mehek (mehek14):

because this one I already did so I knew what to do

Mehek (mehek14):

for 5 you use the functions to explain

OpenStudy (horsegirl27):

ok but idk how to explain that very well

Mehek (mehek14):

for 6, you use \(\Large{f(b) - f(a)/ b - a}\)

Mehek (mehek14):

for 5 just say that the investor's plan has a constant rate

OpenStudy (horsegirl27):

Thank you

Mehek (mehek14):

and the CD's are exponential functions because they are both multiplied by a ratio.

OpenStudy (horsegirl27):

and ik how to do 6 I just hate having to do it

Mehek (mehek14):

ok now 7